# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ I.2 Phonon Dynamical Structure Factor
Experimentally, the direct physical measurable quantity is the phonon dynamical structure factor, $S(\bm{Q},\omega)$, which is determined by the charge density correlation function Hansen and Klein (1976):
$S(\bm{Q},\omega)=\frac{1}{2\pi\hbar}\int_{-\infty}^{\infty}e^{-i\omega t}\langle\rho(\bm{Q},0)\rho^{\dagger}(\bm{Q},t)\rangle\leavevmode\nobreak\ \mathrm{d}t,$ (10)
@@
$\rho(\bm{Q},t)=\sum_{l}\sum_{i}f_{i}(\bm{Q})e^{-i\bm{Q}\cdot[\bm{R}_{l}+\bm{r}_{i}+\bm{u}_{l,i}(t)]},$ (12)
where $l$ represents the primitive cell index, and $i$ labels different atoms within the unit cell. The equilibrium position is given by $\bm{x}_{0}=\bm{R}_{l}+\bm{r}_{i}$ and $\bm{u}_{l,i}(t)$ is the atomic displacement. The dynamical structure factor can then be written as:
$\begin{split}S(\bm{Q},\omega)=\frac{N_{l}}{2\pi\hbar}\sum_{l}e^{i\bm{Q}\cdot\bm{R}_{l}}\sum_{i,j}f_{i}(\bm{Q})f^{*}_{j}(q)e^{-i\bm{Q}\cdot(\bm{r}_{i}-\bm{r}_{j})}\\ \int_{-\infty}^{\infty}e^{-i\omega t}\langle e^{-i\bm{q}\cdot\bm{u}_{0,j}(0)}e^{i\bm{q}\cdot\bm{u}_{l,i}(t)}\rangle\leavevmode\nobreak\ \mathrm{d}t.\end{split}$ (13)
@@
$\begin{split}\langle e^{-i\bm{Q}\cdot\bm{u}_{l^{\prime},j}(0)}e^{i\bm{Q}\cdot\bm{u}_{l,i}(t)}\rangle=e^{-\frac{1}{2}\langle|\bm{Q}\cdot\bm{u}_{l^{\prime},j}|^{2}\rangle}e^{-\frac{1}{2}\langle|\bm{Q}\cdot\bm{u}_{l,i}|^{2}\rangle}\\ \cdot e^{\langle(\bm{Q}\cdot\bm{u}_{l^{\prime}j}(0))(\bm{Q}\cdot\bm{u}_{li}(t))\rangle}.\end{split}$ (14)
We introduce the Debye-Waller factor, $W_{l}\equiv\frac{1}{2}\langle|\bm{Q}\cdot\bm{u}_{li}|^{2}\rangle\approx\frac{3\hbar^{2}|\bm{Q}|^{2}}{2M\beta(k_{B}\Theta_{D})^{2}}$, where $\Theta_{D}$ is the Debye temperature and $\beta=1/k_{B}T$, and define $U\equiv(\bm{Q}\cdot\bm{u}_{l^{\prime},j}(0))$, $V\equiv(\bm{Q}\cdot\bm{u}_{l,i}(t))$. We then obtain:
$\begin{split}\langle e^{-i\bm{Q}\cdot\bm{u}_{l^{\prime},j}(0)}e^{i\bm{Q}\cdot\bm{u}_{li}(t)}\rangle&=e^{-W_{l}}e^{-W_{l^{\prime}}}\\ &(1+\langle UV\rangle+\langle UV\rangle^{2}+...).\end{split}$ (15)
@@
$S(\bm{Q},\omega)_{0p}=N_{l}^{2}\left|\sum_{l}f_{i}(\bm{Q})e^{-W_{d}}e^{i\bm{Q\cdot r_{i}}}\right|^{2}\delta_{\bm{Q}=\bm{G}}\delta(\hbar\omega),$ (17)
$\begin{split}S(\bm{Q},\omega)_{1p}=N_{l}\sum_{\bm{q},\sigma}\frac{1}{\omega_{\sigma}(\bm{q})}\left|\sum_{i}\frac{f_{i}(\bm{Q})}{\sqrt{2M_{i}}}e^{-W_{d}}\bm{Q}\cdot\bm{e}_{\bm{q}\sigma}^{i}e^{i\bm{Q}\cdot\bm{r}_{i}}\right|^{2}\\ \delta_{\bm{Q}-\bm{q}=\bm{G}}\left[\langle n_{\bm{q}\sigma}+1\rangle\delta(\omega-\omega_{\bm{q}\sigma}+\langle n_{\bm{q}\sigma}+1\rangle\delta(\omega+\omega_{\bm{q}\sigma})\right],\end{split}$ (18)
where $\bm{G}$ is the reciprocal lattice vector and $\bm{q}$ is the reduced wave vector in the first Brillouin zone. Using Eq. (18), $S(\bm{Q},\omega)$ can be calculated from first-principles calculations to compare directly with various experimental probes described below. Therefore, it is possible to extract not only the eigenvalues of the dynamical matrix, which correspond to the phonon frequencies, but also the eigenvectors, which are directly related to the intensity of the scattering function $S(\bm{Q},\omega)$.
@@ I.3 Experimental Probes
Phonons can be probed, directly or indirectly, by absorption and scattering techniques. In this section, we provide a concise overview of spectroscopic techniques. For readers seeking in-depth information on instrumental setups and theoretical foundations, we recommend referring to specialized reviews and literature focused on specific technique.
I.3.1 Raman and Infrared Spectrascopy
@@
Raman and infrared (IR) spectroscopy are two well-known methods to probe lattice dynamics with nearly zero momentum transfer Basov and Timusk (2005); Devereaux and Hackl (2007). As we shall see, these techniques have opposite and complementary selection rules, and must be used in combination to measure all phonon modes near the Brilouin zone center.
@@
Figure 2: Schematics of photon scattering processes. (a) depicts the IR spectroscopy. When the incident photon energy, $\hbar\omega$, matches the phonon excitation energy, $E_{ph}$, the incident light will be absorbed, resulting reduced reflection or transmission light intensity. (b) shows the Raman scattering process. The incident photon drives the initial state to a virtual excited state. The system then relax to a final state and emit a photon. Following the energy and momentum conservation, $E_{ph}=\hbar\omega-\hbar\omega^{\prime}$ and $\bm{P}_{ph}=\hbar\bm{k}-\hbar\bm{k}^{\prime}$, where $\hbar\bm{k}$ and $\hbar\bm{k}^{\prime}$ are incident and scattered photon momentum. (c) illustrates the direct RIXS process. Similar to the Raman scattering, RIXS is also a scattering process. The incident photon energy is tuned to a element specific resonant energy. The core electron is excited to an intermediate state. The core hole is then filled by an electron from the conduction band.
@@
The physical principle underlying IR spectroscopy is shown in Fig. 2(a). This technique involves measuring the absorption or reflection of IR radiation. Strong IR absorption occurs when the charge oscillations induced by the optical field resonate with lattice vibrations. Since the electrical dipole moment generated by these charge oscillations is parity-odd, only phonon modes with odd paity are IR-active.
@@
In contrast, Raman scattering is a “photon-in-photon-out” scattering process, as depicted in Fig. 2(b). It measures the the change in wavelength of light as a phonon is either created or annihilated Devereaux and Hackl (2007). Raman scattering involves a change in the polarizability of the crystals. The light-induced dipole moment is:
$\bm{u}=\tilde{\alpha}\cdot\bm{E}_{i}\cos(\bm{k}_{0}\cdot\bm{r}-\omega_{0}t),$ (19)
@@
where $\tilde{\alpha}$, a second rand symmetric tensor, is the Raman polarizability tensor. Since the inelastic scattering of the incident light $\bm{E}_{i}$ can excite atomic vibrations, the polarizability tensor has frequency-dependent contributions at the vibration frequencies:
$\tilde{\alpha}=\tilde{\alpha}_{0}+\sum_{\sigma,\bm{q}}X_{\bm{q}\sigma}\tilde{\alpha}_{\bm{q}\sigma}\cos(\bm{q}\cdot\bm{r}-\omega_{\bm{q}\sigma}t).$ (20)
@@
$\begin{split}\bm{u}=\tilde{\alpha}_{0}\cdot\bm{E}_{i}\cos(\bm{k}_{0}\cdot\bm{r}-\omega_{0}t)+\\ \sum_{\bm{q},\sigma}X_{\bm{q},\sigma}\tilde{\alpha}_{\bm{q}\sigma}\cdot\bm{E}_{i}\cos[(\bm{k}_{0}\pm\bm{q})\cdot\bm{r}-(\omega\pm\omega_{\bm{q}\sigma})t].\end{split}$ (21)
@@
In Eq. (21), the $+\omega$ and $-\omega$ photon frequency shifts are known as Stokes and anti-Stokes process in the Raman scattering, respectively. For centrosymmetric materials, since both $\bm{u}$ and $\bm{E}$ are parity-odd, $\tilde{\alpha}_{\sigma}$ must be parity-even to ensure that the Raman-active modes are symmetry-allowed.
I.3.2 Inelastic x-ray and Neutron Scattering
@@
Inelastic x-ray (IXS) and inelastic neutron scattering (INS) are powerful techniques to directly measure the dynamic structure factor $S(\bm{Q},\omega)$ of the bulk phonons across a broad range of energies and momenta Feigin et al. (1987); Higgins and Maconnachie (1986); Kotani and Shin (2001). To begin with, we first introduce the interaction between the photon field and charged particles:
$\mathcal{H}_{\textrm{I}}=\frac{e\bm{A\cdot P}}{m}+\frac{e^{2}A^{2}}{2m}.$ (22)
The vector potential $\bm{A}$ can be written in photon operators, $b_{u\bm{k}}$ and $b^{\dagger}_{u\bm{k}}$, as:
$\bm{A}(\bm{r},t)=\sum_{u}\sum_{\bm{k}}\hat{\varepsilon}_{u}\sqrt{\frac{\hbar}{2\epsilon_{0}V\omega_{\bm{k}}}}[b_{u\bm{k}}e^{i(\bm{k\cdot r}-\omega t)}+b^{\dagger}_{u\bm{k}}e^{-i(\bm{k\cdot r}-\omega t)}],$ (23)
where $\hat{\varepsilon}_{u}$ represents the polarization of the light. In the case of IR spectroscopy, the cross-section is determined by the first term on the right-hand side of Eq. (22), which corresponds to single-photon absorption processes. For Raman scattering, both term contribute to the cross-section. In contrast, for (non-resonant) IXS, the second term dominates the cross-section. The Thomson differential cross-section, for a solid angle $\mathrm{d}\Omega$ and energy bandwidth $\mathrm{d}\omega$, is given by:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\frac{k_{f}}{k_{i}}r_{0}^{2}|\hat{\varepsilon}_{i}\cdot\hat{\varepsilon}_{f}|^{2}S(\bm{Q},\omega),$ (24)
where $i$ and $f$ denote the initial and final states. $r_{0}=\frac{e^{2}}{mc^{2}}\sim$2.81 femtometer (fm) is the classical radius of the electron. For a typical IXS experiment, the energy transfer, $\omega$, is about 5$\sim$6 orders magnitude smaller than the incident photon energy; therefore, $\frac{k_{f}}{k_{i}}\sim$1, and $\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}\propto S(\bm{Q},\omega)$.
@@
For INS, incident neutrons interact with atomic nuclei through short-range nuclear forces, which operate within the femtometer length scale. The coherent scattering contribution to the INS cross-section of INS is given by:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\frac{\sigma_{\textrm{coh}}}{4\pi}\frac{k_{f}}{k_{i}}NS(\bm{Q},\omega).$ (25)
@@
While both IXS and INS are powerful techniques for probing phonon dynamics, they are often complementary in many aspects. For instance, the scattering intensity in IXS approximately scales with $Z^{2}$, where $Z$ is the atomic number, whereas INS displays complex isotope effect. As a result, the choice between the two techniques may depend on the specific chemical elements present in the sample.
@@
In addition, IXS offers higher momentum resolution and the ability to study small-size samples ($\sim 10\leavevmode\nobreak\ \mu m\times 30\leavevmode\nobreak\ \mu m$) due to the high brilliance of modern synchrony light source. On the other hand, INS can achieve higher energy resolution and enables the simultaneous collection of data across a large momentum space using time-of-flight spectrometer.
I.3.3 Resonant Inelastic x-ray Scattering

### v2

@@ I.2 Phonon Dynamical Structure Factor
Experimentally, the direct physical measurable quantity is the phonon dynamical structure factor, $S(\bm{Q},\omega)$, which is determined by the charge density correlation function hansen1976dynamical :
$S(\bm{Q},\omega)=\frac{1}{2\pi\hbar}\int_{-\infty}^{\infty}e^{-i\omega t}\langle\rho(\bm{Q},0)\rho^{\dagger}(\bm{Q},t)\rangle\penalty 10000\ \mathrm{d}t,$ (12)
@@
$\rho(\bm{Q},t)=\sum_{i}\sum_{s}f_{s}(\bm{Q})e^{-i\bm{Q}\cdot[\bm{R}_{i}+\bm{r}_{s}+\bm{u}_{i,s}(t)]},$ (14)
where $i$ represents the primitive cell index, and $s$ labels different atoms within the unit cell. The equilibrium position is given by $\bm{x}_{0}=\bm{R}_{i}+\bm{r}_{s}$ and $\bm{u}_{i,s}(t)$ is the time-dependent atomic displacement. The dynamical structure factor can then be written as:
$\begin{split}S(\bm{Q},\omega)=\frac{N_{l}}{2\pi\hbar}\sum_{l}e^{i\bm{Q}\cdot\bm{R}_{l}}\sum_{s,s^{\prime}}f_{s}(\bm{Q})f^{*}_{s^{\prime}}(q)e^{-i\bm{Q}\cdot(\bm{r}_{s}-\bm{r}_{s^{\prime}})}\\ \int_{-\infty}^{\infty}e^{-i\omega t}\langle e^{-i\bm{q}\cdot\bm{u}_{0,s^{\prime}}(0)}e^{i\bm{q}\cdot\bm{u}_{l,s}(t)}\rangle\penalty 10000\ \mathrm{d}t.\end{split}$ (15)
@@
$\begin{split}\langle e^{-i\bm{Q}\cdot\bm{u}_{l^{\prime},s^{\prime}}(0)}e^{i\bm{Q}\cdot\bm{u}_{l,s}(t)}\rangle=e^{-\frac{1}{2}\langle|\bm{Q}\cdot\bm{u}_{l^{\prime},s^{\prime}}|^{2}\rangle}e^{-\frac{1}{2}\langle|\bm{Q}\cdot\bm{u}_{l,s}|^{2}\rangle}\\ \cdot e^{\langle(\bm{Q}\cdot\bm{u}_{l^{\prime}s^{\prime}}(0))(\bm{Q}\cdot\bm{u}_{ls}(t))\rangle}.\end{split}$ (16)
We introduce the Debye-Waller factor, $W_{l}\equiv\frac{1}{2}\langle|\bm{Q}\cdot\bm{u}_{ls}|^{2}\rangle\approx\frac{3\hbar^{2}|\bm{Q}|^{2}}{2M\beta(k_{B}\Theta_{D})^{2}}$, where $\Theta_{D}$ is the Debye temperature and $\beta=1/k_{B}T$, and define $U\equiv(\bm{Q}\cdot\bm{u}_{l^{\prime},s^{\prime}}(0))$, $V\equiv(\bm{Q}\cdot\bm{u}_{l,s}(t))$. We then obtain:
[…]
@@
$S(\bm{Q},\omega)_{0p}=N_{l}^{2}\left|\sum_{s}f_{s}(\bm{Q})e^{-W_{d}}e^{i\bm{Q\cdot r_{s}}}\right|^{2}\delta_{\bm{Q}=\bm{G}}\delta(\hbar\omega),$ (19)
$\begin{split}S(\bm{Q},\omega)_{1p}=N_{l}\sum_{\bm{q},\sigma}\frac{1}{\omega_{\sigma}(\bm{q})}\left|\sum_{s}\frac{f_{s}(\bm{Q})}{\sqrt{2M_{s}}}e^{-W_{d}}\bm{Q}\cdot\bm{\epsilon}_{\bm{q}\sigma}(s)e^{i\bm{Q}\cdot\bm{r}_{s}}\right|^{2}\\ \delta_{\bm{Q}-\bm{q}=\bm{G}}\left[\langle n_{\bm{q}\sigma}+1\rangle\delta(\omega-\omega_{\bm{q}\sigma}+\langle n_{\bm{q}\sigma}+1\rangle\delta(\omega+\omega_{\bm{q}\sigma})\right],\end{split}$ (20)
where $\bm{G}$ is the reciprocal lattice vector and $\bm{q}$ is the reduced wave vector in the first Brillouin zone. Using Eq. (20), $S(\bm{Q},\omega)$ can be calculated from first-principles calculations to compare directly with various experimental probes described below. By measuring $S(\mathbf{Q},\omega)$, one can extract not only the eigenvalues of the dynamical matrix, which correspond to the phonon frequencies, but also the eigenvectors, which are directly related to the scattering intensity.
@@ I.3 Experimental Probes
Phonons can be probed, directly or indirectly, by absorption and scattering techniques. In this section, we provide a concise overview of spectroscopic techniques. For readers seeking in-depth information on a specific experimental probe, we refer to specialized technical review articles.
I.3.1 Light-matter interactions
For light-matter interactions, the scattering and absorption cross-sections are related to the transition, $W$, between initial state, $|i\rangle$, and final state, $|f\rangle$, that is give by Fermi’s golden rule:
$W=\frac{1}{\mathcal{Z}}\sum_{i,f}e^{-\beta\epsilon_{i}}|M_{if}|^{2}\delta(\epsilon_{f}-\epsilon_{i}-\hbar\Omega)$ (21)
$\mathcal{Z}$ is the partition function. $M_{if}=\langle f|H_{I}|i\rangle$ is the scattering matrix element, where $H_{I}$ is the scattering operator describing the interactions between the photon field and charged particles:
$\mathcal{H}_{\textrm{I}}=\frac{e\bm{A\cdot P}}{m}+\frac{e^{2}A^{2}}{2m}.$ (22)
The vector potential $\bm{A}$ can be written in photon operators, $b_{u\bm{k}}$ and $b^{\dagger}_{u\bm{k}}$:
$\begin{split}\bm{A}(\bm{r},t)=\sum_{u}\sum_{\bm{k}}\hat{\varepsilon}_{u}\sqrt{\frac{\hbar}{2\epsilon_{0}V\omega_{\bm{k}}}}[b_{u\bm{k}}e^{i(\bm{k\cdot r}-\omega t)}\\ +b^{\dagger}_{u\bm{k}}e^{-i(\bm{k\cdot r}-\omega t)}],\end{split}$ (23)
where $\hat{\varepsilon}_{u}$ is the photon polarization vector. A general expression for the differential light scattering cross section is:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\hbar r_{0}^{2}\frac{\omega_{f}}{\omega_{i}}W,$ (24)
where $r_{0}=\frac{e^{2}}{mc^{2}}\sim$2.81 femtometer (fm) is the Thompson radius of the electron. The $\omega_{i}$ and $\omega_{f}$ represent the energy of incident and scattered photons with momentum $\mathbf{k}_{i}$ and $\mathbf{k}_{f}$.
I.3.2 Raman and Infrared Spectroscopy
@@
Raman and infrared (IR) spectroscopy are well-known methods to probe lattice dynamics with nearly zero momentum transfer Basov2005Electrodynamics ; Devereaux2007inelastic . In the presence of inversion symmetry, these techniques have opposite and complementary selection rules, and must be used in combination to measure all phonon modes near the Brillouin zone center.
@@
Figure 2: Schematics of photon scattering processes. (a) depicts the IR spectroscopy. When the incident photon energy, $\hbar\omega$, matches the phonon excitation energy, $E_{ph}$, the incident light will be absorbed, resulting reduced reflection or transmission light intensity. (b) shows the Raman scattering process. The incident photon drives the initial state to a virtual excited state. The system then relax to a final state and emit a photon. Following the energy and momentum conservation, $E_{ph}=\hbar\omega-\hbar\omega^{\prime}$ and $\bm{Q}_{ph}=\hbar\bm{k}-\hbar\bm{k}^{\prime}$, where $\hbar\bm{k}$ and $\hbar\bm{k}^{\prime}$ are incident and scattered photon momentum. (c) illustrates the direct RIXS process. Similar to the Raman scattering, RIXS is also a scattering process. The incident photon energy is tuned to a element specific resonant energy. The core electron is excited to an intermediate state. The core hole is then filled by an electron from the conduction band.
@@
The physical principle underlying IR spectroscopy is shown in Fig. 2(a). This technique involves measuring the absorption or reflection of IR radiation. Strong IR absorption occurs when the charge oscillations induced by the optical field resonate with lattice vibrations. In this process, the initial and final states contains one and zero photon, respectively. The cross-section is therefore determined by the parity-odd dipole term in Eq. 22, $\bm{A\cdot P}$. Since the zero-phonon initial state is parity-even, the final states must be parity-odd phonon modes.
@@
In contrast, Raman scattering is a “photon-in-photon-out” scattering process, as depicted in Fig. 2(b). It measures the the change in wavelength of light as a phonon is either created or annihilated Devereaux2007inelastic . In this process, it measures the polarizability of the crystals under the photon field, where both linear and quadratic terms in Eq. 22 contribute to the cross-section. To understanding the selection rule, we can consider the light-induced dipole moment:
$\bm{D}=\tilde{\alpha}\cdot\bm{E}_{i}\cos(\bm{k}_{0}\cdot\bm{r}-\omega_{0}t),$ (25)
@@
where $\tilde{\alpha}$, a second-rank symmetric tensor, is the Raman polarizability tensor. Since the inelastic scattering of the incident light $\bm{E}_{i}$ can excite atomic vibrations, the polarizability tensor has frequency-dependent contributions at the vibration frequencies:
$\tilde{\alpha}=\tilde{\alpha}_{0}+\sum_{\sigma,\bm{q}}\tilde{\alpha}_{\bm{q}\sigma}\cos(\bm{q}\cdot\bm{r}-\omega_{\bm{q}\sigma}t).$ (26)
@@
$\begin{split}\bm{D}=\tilde{\alpha}_{0}\cdot\bm{E}_{i}\cos(\bm{k}_{0}\cdot\bm{r}-\omega_{0}t)+\\ \sum_{\bm{q},\sigma}\tilde{\alpha}_{\bm{q}\sigma}\cdot\bm{E}_{i}\cos[(\bm{k}_{0}\pm\bm{q})\cdot\bm{r}-(\omega\pm\omega_{\bm{q}\sigma})t].\end{split}$ (27)
@@
In Eq. (27), the $\omega$ and $\omega$ photon frequency shifts are known as Stokes and anti-Stokes process in the Raman scattering, respectively. For centrosymmetric materials, since both $\bm{D}$ and $\bm{E}$ are parity-odd, $\tilde{\alpha}_{\sigma}$ must be parity-even to ensure that the Raman-active modes are symmetry-allowed.
Beyond this simply selection rule, since Raman scattering involves virtual intermediate states, its scattering matrix element, $M_{if}$, display strong photon polarization dependence Devereaux2007inelastic . By controlling the photon polarization, $\hat{\epsilon}_{u}$, Raman can selectively probe phonon modes with different chirality ishito2023truly .
I.3.3 Inelastic x-ray and Neutron Scattering
@@
Inelastic x-ray (IXS) and inelastic neutron scattering (INS) are powerful techniques to directly measure the dynamic structure factor $S(\bm{Q},\omega)$ of bulk phonons across a broad energy and momentum range feigin1987structure ; higgins1986neutron ; kotani2001resonant .
For (non-resonant) IXS, the scattering cross-section is determined by the quadratic in $\mathbf{A}$ term in Eq. 22. The Thomson differential cross-section is given by:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\frac{k_{f}}{k_{i}}r_{0}^{2}|\hat{\varepsilon}_{i}\cdot\hat{\varepsilon}_{f}|^{2}S(\bm{Q},\omega).$ (28)
[…]
@@
Different from x-ray scattering, INS arises from the energy and momentum transfer between incident neutrons and the atomic nuclei. These interactions are short-range nuclear forces, which operate within the femtometer length scale. This length-scale is much smaller than the wavelength of neutrons. For this reason, the scattering center can be viewed as a delta function, corresponding to momentum independent form factor that is different from photon probes. The coherent scattering contribution to the INS cross-section of is given by:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\frac{\sigma_{\textrm{coh}}}{4\pi}\frac{k_{f}}{k_{i}}NS^{N}(\bm{Q},\omega),$ (29)
where $S^{N}(\bm{Q},\omega)$ is similar to $S(\bm{Q},\omega)$ except $f_{i}(\mathbf{Q})=1$ for the neutron scattering. The $\sigma_{\textrm{coh}}$ depends on the nuclear isotope, spin relative to the neutron, and nuclear eigenstate.
@@
While both IXS and INS are powerful techniques for probing phonon dynamics, they are often complementary in many aspects. For instance, the scattering intensity of IXS approximately scales with $Z^{2}$, where $Z$ is the atomic number, whereas INS displays complex isotope effect. As a result, the choice between the two techniques may depend on the specific chemical elements present in the sample.
@@
In addition, IXS offers higher momentum resolution and the ability to study small-size samples ($\sim 10\penalty 10000\ \mu m\times 30\penalty 10000\ \mu m$) due to the high brilliance of modern synchrony light source. On the other hand, INS can achieve higher energy resolution and allow simultaneous collection of data across a large momentum space using time-of-flight spectrometer.
I.3.4 Resonant Inelastic x-ray Scattering